Site C0002

RWD was employed for the first time by IODP during Expedition 338 in order to allow cutting the 12¼ inch diameter pilot hole and opening the hole to 20 inches at the same time. This procedure was employed to facilitate the installation of casing strings; however, these strings were not installed because of the early termination of riser operations (see “Site C0002” operations below). There is a concentric hole opener between the bit and the underreamer. The underreamer used to enlarge the hole to 20 inches was the National Oilwell Varco Anderreamer. The design of the bottom-hole assembly (BHA) also included a complete LWD tool suite. During riserless coring, no underreamer was used and the bit had a 12¼ inch diameter and used the standard RCB coring system for Holes C0002H, C0002I, and C0002J. Holes C0002K and C0002L were cored with a HPCS, EPCS, and ESCS (Table T2).

Site C0012

Hole C0012H was drilled with an LWD tool string similar to that used for logging in Hole C0002F; however, no underreamer was used. The 12¼ inch polycrystalline diamond compact (PDC) bit was employed to allow drilling in soft and semi-indurated sediment and into basement.

Site C0018

Hole C0018B was drilled with an LWD tool string similar to that used for logging in Hole C0002F; however, neither the underreamer nor the sonicVISION tool were used. The 12¼ inch PDC bit was employed to allow drilling in soft and semi-indurated sediment and into basement.

Site C0021

Hole C0021A was drilled with the same LWD tool string that was used for logging in Hole C0018B. Hole C0021B was cored with HPCS and EPCS.

Site C0022

Hole C0022A was drilled with the same LWD tool string that was used for logging in Hole C0018B. Hole C0021B was cored with HPCS, EPCS, and ESCS.

Shimizu, Japan, port call

Expedition 338 officially began at 0000 h on 1 October 2012 while the Chikyu was north of Izu-Oshima, Japan, as part of the evacuation procedures for Typhoon Jelawat. The evacuation ended at 0400 h on 1 October, when the vessel began to return to port call in Shimizu, Japan. Port call began on 0550 h on 2 October, with all loading operations concluded by 2400 h on 3 October. The vessel departed Shimizu, Japan, for Site C0002 at 1035 h on 4 October.

Site C0002

Seabed survey and transponder deployment began at 1230 h on 5 October and were completed on 7 October 2012. After calibrating transponders, the Chikyu sailed to the BOP and riser running point 20.8 nmi northwest of Site C0002. A BOP and riser joints were set up in the moonpool and prepared for running beginning at 0400 h on 8 October. By 0400 h on 9 October, the BOP reached 490 m drilling depth below rig floor (DRF); however, because of the development of Typhoon Prapiroon, a decision was made to wait on weather (WOW) and follow the typhoon track at 0900 h on 9 October; subsequently, the BOP and riser began to be recovered for evacuation standby at 1300 h on 11 October, and recovery was completed at 0600 h on 12 October. The vessel moved 21 nmi northwest of Site C0002 to facilitate the arrival of the first scientists by helicopter transfer on 13 October and to avoid the strengthening current. The Chikyu continued to observe developing typhoon tracks, remaining in WOW status until 0700 h on 19 October, when preparations to begin running the BOP into the moonpool commenced. Once the BOP was in the moonpool at ~1700 h, BOP pressure tests were run, ending at 1800 h when the BOP was run into the water. Careful attention was paid to the auxiliary (AUX) line and buoyant riser joints, all the while conducting a schedule of regular AUX pressure tests (six pressure tests in total). AUX line troubleshooting and buoyancy riser joint replacement was required on 21 October when one joint was replaced.

While running the BOP, several function tests and pressure tests found failures in the AUX line, conduit line, and hot line, which were subsequently repaired. The BOP finally landed on the wellhead at 2130 h on 26 October, and all function tests, repairs, and maintenance were complete by 1045 h on 29 October. The tests included a pickup test by increasing tensioner tension, slump tests by decreasing drawworks tension, function tests of the diverter, pressure tests of the wellhead connector and the 20 inch casing, function tests of the BOP from both the Blue Pod (driller’s control panel) and the Yellow Pod (toolpusher’s control panel), a pressure test of the BOP, a pressure test of the BOP with 20 inch casing and inside drill pipe, a function test of the remaining valves in block position, and confirmation of working time and flow rate of the BOP. The BHA for drill-out cement (DOC) was run into the hole at 1900 h on 30 October and tagged the bottom (842 mbsf) at 1415 h on 31 October. DOC with the 17 inch bit reached 872.5 mbsf, with an extension into the formation (to 875.5 mbsf) to confirm the cement layer was completely drilled through ended at 2045 h. After drilling was complete, the hole was swept with Hi-Vis mud, and then seawater, followed by KNPP mud. Methane (15.5%) was found in the drilling mud, and after monitoring, it was decided to move ahead with two LOTs of the 20 inch casing shoe starting at ~1100 h on 1 November. The DOC BHA was recovered and laid down, after which the LWD/MWD underreamer BHA was made up and run into the hole (Table T3) and tested at 0430 h on 2 November. All function tests (n = 4: shallow; 50 meters below sea level [mbsl]; 132.5 mbsf; and 778.5 mbsf) were successfully completed by 0100 h on 3 November, after which initial pilot hole drilling began. The BHA was picked up after drilling ahead to 915.19 mbsf, back to 905.8 mbsf at 0745 h on 4 November to activate the underreamer (total of six attempts made), which took ~2 h.

Drilling with 12¼ × 20 inch LWD/MWD with underreamer began at 1000 h on 4 November. Rate of penetration (ROP) was controlled at 40 m/h from 914.5 mbsf and changed to 20–26 m/h below 923.8 mbsf. Concerns with the large volume of cuttings, more than the waste mud control system could cope with, required careful control, sweeping, and sometimes suspension of drilling. Repeat logging was carried out three times using these periods (~30 m/h): uplogging from 1495.5 to 1457.5 mbsf, downlogging from 1503.5 to 1538.5 mbsf, and downlogging from 1557.5 to 1615.5 mbsf. Expected rough weather caused drilling to be suspended at 1538.5 mbsf (2200 h on 5 November), and the BHA was pulled out of the hole above the BOP. Drilling resumed (ROP = 15 m/h) at 2100 h on November 6 but stopped again when 1604.5 mbsf was reached because black smoke flow was observed coming from one port on the BOP 36 inch pressure housing with subsequent hydrate accumulation around the wellhead (1230 h on 7 November). Observation for 6 h showed stable flow, so drilling was resumed (ROP = 10–15 m/h) at 2300 h on 7 November. From 1604.5 to 2005.5 mbsf (2300 h on 7 November to 0800 h on 11 November), drilling continued with some interruptions, mainly because of waste mud control system issues (e.g., mud pump problems, screw conveyer for mud transfer issues, and strong currents interfering with supply boats). Drilling was stopped again to standby for rough weather, and the BHA was pulled out of the hole above the BOP at 1515 h on 12 November.

During WOW, a critical function failure of the Double V Shear (DVS) ram (a key component of the BOP designed to close the hole while maintaining annular pressure after the BOP disconnects) was found. It was decided at 0530 h on 13 November to suspend the hole until the failure was resolved. Fortunately, the root cause of the failure was determined, and a mitigation plan was enacted for the BOP; therefore, at 0230 h on 14 November, the BOP disconnection plan was canceled. DOC began at 0115 h on 15 November and completed at 1630 h when the DOC BHA was pulled out of the hole to the surface. Subsequently, the Anderreamer BHA for reaming and enlarging the hole from 12¼ to 20 inches was run into the hole at 0000 h on 16 November. The Anderreamer was activated at 2848 m DRF at 0300 h; reaming was conducted from 0300 to 2300 h on 16 November. When bit depth reached 3955 m DRF (1990 mbsf) and the Anderreamer depth and 20 inch hole was at 3949 m DRF (1984 mbsf), an approaching weather system caused another WOW. Reaming and backreaming down to the final bit depth were performed over the next 5 h until the BHA was pulled out of the hole to 1576 m DRF by 1430 h on 17 November after performing backreaming in the hole and spotting high-viscosity mud water. As a standby operation, the kill line was flushed and the DVS ram was closed.

At 1730 h on 17 November, the vessel went into emergency disconnect as the current speed of ~4.5 kt from the west suddenly changed direction along with strong winds (~27 m/s), the combination of which forced the vessel 40 m from the well center, and control of the vessel was lost. The emergency disconnect was conducted both safely and efficiently as the vessel drifted 1300 m east of the wellhead at 4 kt. However, while this was happening, the riser intermediate flex joints struck the hull of the vessel and sustained damage. Pulling out of the hole to the surface continued, ending at 0415 h on 18 November; whereupon the vessel recovered the ROV and moved to the low-current area (LCA) 13 nmi northwest of Hole C0002F to begin riser recovery. The tensioner ring was removed, and electrical cable connections were removed as well. Riser pullout began at 2300 h and termination joints were checked. The diverter was laid out from 0000 h on 19 November, and the riser termination joint was pulled out of the moonpool at 0430 h. The hot line was reterminated in the moonpool, and a pressure check from 1400 to 1800 h confirmed the BOP status was still good. The gooseneck was inspected, including welding points on the slip joint with magnetic particle inspection (MPI) for cracks from 2100 to 0030 h on 20 November. Riser sections continued to be pulled, stopping to perform pressure tests on the choke and kill lines for 3 hours from 0315 h. After the intermediate flex joint was pulled up, the termination joint weld was inspected via MPI, whereupon four cracks were found; advice on repair was sought from the manufacturer. The results of careful inspection and manufacturer advice led to the conclusion that the repairs needed could only be conducted on shore; accordingly, it was decided to end riser operations once it was clear that the time needed for repairs would exceed the total planned expedition duration. Meanwhile, a pressure test of the conduit and kill lines on the gooseneck passed inspection after several tests and leakage troubleshooting by 2315 h on 21 November. Once passed, the gooseneck was removed so the intermediate flex joint could be installed, finishing at 0600 h on 22 November. Once installed, the gooseneck and choke, booster, and kill moonpool hoses were made up, pressure-tested, and installed by 1515 h. The landing and riser joints were also picked up, connected, and lowered to the moonpool, where work began to couple the riser tensioners to the landing ring, finishing by 0330 h on 23 November. After the ROV was launched, a dummy-landing test 50 m from the well center was conducted successfully at 1730 h, and by 2030 h the lower marine riser package (LMRP) was landed on the lower BOP stack and locked in place. Riser running components were laid down by 0600 h on 24 November. The diverter assembly was made up and run into the hole to 2987 m DRF by 1700 h, and then was slowly pulled out of the hole to 2878 m DRF. The first cement plug was set at 2200 h, and then pulled out of the hole to 2460 m DRF while waiting for cement. At 0445 h on 25 November, the BHA was run into the hole back to 2805 m DRF to confirm the top of cement and begin the first cement plug pressure test. After confirming the cement plug, the BHA was pulled out of the hole to 2367 m DRF to set the second cement plug, pulled out of the hole to 2120 m DRF, and then pulled out of the hole to the surface by 1745 h. The Hydralift Power Swivel (HPS) was parked and the riser running tool and riser guide head were installed, finishing at 2245 h. Once complete, the master bushing was removed while the ROV removed the hydrate build-up around the wellhead connector. The diverter and upper flex joint were picked up and laid down by 0400 h on 26 November, whereupon the BOP was disconnected, the wellhead was examined by the ROV, and the ROV was recovered to the surface by 0730 h. The vessel moved to the LCA ~15 nmi from Hole C0002F and began recovering riser joints and the BOP at 1200 h on 27 November. The intermediate flex joint was recovered at 0200 h on 28 November, after which riser joint and BOP recovery continued. Recovery was briefly halted on 30 November to load the guide horn from the supply boat Shincho-Maru and resumed at 1000 h. The riser joints were all recovered and laid down by 0930 h on 1 December, after which the BOP was pulled out to the surface, landed on the BOP cart, and moved to its storage position aft of the moonpool by 1815 h. Once the BOP was loaded on the cart, the HPS was rigged up again, and the vessel returned to Site C0002, arriving at 0630 h on 1 December. At Site C0002, the ROV dove at 0745 h, and the vessel shifted to the Hole C0002G long-term borehole monitoring system (LTBMS) observatory for ROV inspection. The vessel shifted back to Hole C0002F, and the ROV set the corrosion cap and checked the bull’s eyes, finishing by 1145 h. Once complete, the ROV began simultaneously recovering and deploying transponders; 4 transponders were deployed and 10 were recovered, all completed by 0330 h on 2 December. The lower and middle guide horn sections were set on the BOP cart and connected by 0415 h; once connected, the vessel moved to Site C0012.

Hole C0002H

On 8 December at 0130 h, the ship moved 7 nmi north-northwest of Hole C0002H after preparations for coring began with the makeup of the RCB BHA. There was a short WOW lasting until 1600 h, after which the vessel moved to within 5 nmi west of the well center. Drifting in began at 1800 h while the BHA was run into the hole, with a 1 h standby as an internal blowout preventer ball valve malfunction was solved. We continued to drift in to the well center, dropping the center bit at 0300 h on 9 December. The seafloor was tagged and confirmed at 1965 m DRF (1936.5 mbsl) at 0345 h, as indicated by an increase in WOB. We washed down the first 36 m then began drilling a 10⅝ inch hole at 0430 h, reaching 752 mbsf by 2000 h. The sinker bar was run down to recover the center bit for a wear check at 2145 h and dropped again at 2330 h, landing 13 min later. We then drilled down to 1055 mbsf by 0730 h on 10 December. After a series of sweeping out the hole, drilling ahead, and sweeping again, the sinker bar was run at 1100 mbsf to recover the center bit in preparation for dropping the inner core barrel before coring began. At 1430 h, the first coring began, advancing 9.5 m to 1110 mbsf before being recovered on deck at 1545 h. The BHA was stuck for the next 45 min, but after working and sweeping, the core barrel was dropped at 1845 h for the next coring advance. Coring to 1120 mbsf (9.5 m coring advance) began and was recovered on deck at 2105 h. After the core barrel was recovered, the drill pipe was stuck in the hole again, but constant work freed the pipe at 0145 h on 11 December. The BHA was pulled out of the hole with reaming and laid down on deck by 0715 h. No obvious overpull or packoff indication was seen on the drill string, so preparations to return to RCB coring began. The vessel was moved 2 nmi from the next well for preparation for running the RCB BHA into the hole.

Hole C0002I

Hole C0002I operations began with drifting in at 1100 h on 11 December and dropping the center bit at 1630 h prior to spudding in at 1964.5 m DRF (1936 mbsl) 9 min later. The first 33.5 m was washed down and drilling ahead began at 1800 h, reaching 818.5 mbsf at 0600 h on 12 December. Another series of drilling ahead, sweeping, and hole cleaning began once past 905 mbsf. On 13 December, the bit reached 1360 mbsf at 0430 h, continuing until 0445 h on 13 December, when the 4S azimuth thruster shut down and dynamic positioning (DP) status changed to “yellow.” Once this occurred, the BHA was pulled out of the hole to 900 mbsf in preparation for emergency pulling out of the hole to the seafloor. The 4S azimuth thruster was restarted at 0600 h, and DP status returned to “green.” Circulation and hole cleaning began once “green” status was recovered and drilling back to 1105.5 mbsf stopped at 1005 mbsf, whereupon the BHA became stuck at 1030 h on 13 December. Operations to free the stuck pipe began immediately. Attempts to recover the center bit at 1400 h on 14 December were unsuccessful; therefore, rig up of the Schlumberger wireline tool, the Free-Point Indicator Tool (FPIT), began on 1645 h. The stuck position was confirmed at 853 and 915 mbsf, after which the FPIT was rigged down for running the colliding tool to free the drill pipe on 0300 h on 15 December. The colliding tool was set and the explosive primed at 0700 h, when operations to install the colliding tool were suspended because of bad weather conditions, specifically high winds. The lower connection of the drill pipe on the rig floor was broken to run the colliding tool directly from the rig floor by 0745 h when the Schlumberger wireline winch failed. Troubleshooting began immediately, finishing at 2130 h. The colliding tool was rigged up and the explosive was reset by 2315 h. The tool was run at 0300 h on 16 December, and the drill pipe was cut at 0345 h. The colliding tool was rigged down by 0515 h, and pulling out of the hole began. After the tools were laid down, the Chikyu moved to the LCA to load equipment and to perform maintenance on the HPS top drive, ending on 17 December. After pressure tests confirmed the integrity of the HPS repairs, we began making up and testing the RCB BHA for Hole C0002J.

Hole C0002J

The vessel moved to a position 3 nmi west of Hole C0002J, while the RCB BHA was made up and run into the hole from 1830 h on 17 December. Spudding in Hole C0002J was confirmed at 1966 m DRF (1937.5 mbsl) at 0830 h on 18 December, and the BHA immediately jetting in the first 35 mbsf before drilling began. By 2330 h, the bit reached 872.5 mbsf before stopping to space out a single joint. Drilling down continued until reaching 902 mbsf at 0045 h on 19 December, and after sweeping out the hole, coring began from 0430 h. A total of seven RCB cores were collected, finishing at a total depth (TD) of 940 mbsf at 1930 h. Once coring was completed, kill mud was spotted and the RCB BHA was pulled out of the hole to the surface, and tool lay down was completed by 0545 h on 20 December. The vessel moved upstream 2 nmi to Hole C0002K and began preparations for coring in Hole C0002K.

Hole C0002K

Preparations for HPCS/ESCS coring started from 0545 h on 20 December, and the BHA was run into the hole to 1966 m DRF (estimated water depth = 1937.5 mbsl) by 1645 h. The 11 inch BHA washed down the first 30 mbsf by 1700 h then began drilling ahead to 200 mbsf. From 2230 h, coring began with the HPCS, ending after reaching 204.5 mbsf. The low recovery of the last HPCS core (n = 2) caused the drilling program to switch to the ESCS from 204.5 mbsf. The ESCS assembly preparation was completed by 0215 h on 21 December. The inner barrel was dropped 15 min later when ESCS coring began. A total of 11 ESCS cores were collected from 204.5 to 286.5 mbsf and the last core was cut at 2315 h on 21 December. Once coring was completed, the ESCS BHA was pulled out of the hole to 1780 m DRF by 0100 h on 22 December. Two stands of S-150 drill pipe were added, and then the BHA was run down to 1800 m DRF to WOW from 0300 to 0830 h on 22 December. Once the cold front had passed, the vessel was shifted to the Hole C0002L well center for more coring.

Hole C0002L

The ESCS was run down from 1800 to 1960 m DRF by 1030 h, just above the seafloor. Spudding in Hole C0002L began at 1045 h (water depth = 1937.5 mbsl) and proceeded to wash down to 42 mbsf, after which drilling down to 277 mbsf was completed by 1830 h. ESCS coring began at 277 mbsf from 2000 h on 22 December, finishing at a TD of 505.0 mbsf with Core 338-C0002L-24X at 0215 h on 25 December. Once coring operations were completed, 35 m3 of kill mud was spotted in the hole and the BHA was pulled out of the hole to 1932 m DRF (above the seabed) by 0400 h. The BHA was laid down by 1030 h, after which transponders were released and recovered (by 1230 h) by the watch boat, ending Hole C0002L operations. The vessel shifted 2 nmi upstream of Site C0018 in preparation for LWD drilling.

Hole C0012H

The Chikyu arrived at Site C0012 at 1045 h on 2 December; the ROV was prepared with four transponders and dove at 1230 h. The ROV dropped the transponders by 2100 h, after which it was recovered to the surface and the vessel moved to the planned Hole C0012H well center by 2200 h. DP calibration was complete by 2345 h, and the LWD assembly including the sonicVISION tool was made up for running into the hole. The LWD BHA began running into the hole at 0430 h on 3 December. After passing a series of mid-water tests, the LWD tools reached the seafloor when the real-time data signal was lost. After spending some time on troubleshooting operations, the tool was pulled out of the hole to the rig floor for examination and possible replacement, reaching the deck at 0245 h on 4 December. Although a jamming alert from the TeleScope was received and visual inspection found no issues were found with the tool, a spare TeleScope was swapped into the BHA and ran back into the hole at 0645 h on 4 December. A series of tests were again run as the BHA was run down to the seafloor, and once all tests had been completed successfully, the LWD BHA was ready to spud in. Using the LWD tools to determine the mudline, the seafloor was set at 3538 m DRF (3509.5 mbsl) and drilling ahead resumed from 1730 h. The Schlumberger logging engineers were concerned that the sonicVISION configuration should be changed by downlink because in their estimation the real-time sonic velocity was not reliable. After close consultations with the Co-Chief Scientists, it was decided to leave the sonicVISION configuration as is and continue drilling from 3686 m DRF at 2315 h on 4 December. At 0430 h on 6 December, the planned TD of 4238 m DRF (700 mbsf) was reached, after which an additional 10 m was drilled ahead to allow the sonicVISION tool to record an interval where resistivity was seen to increase. The final TD in Hole C0012H of 4248 m DRF (710 mbsf) was reached at 0515 h, when circulation, a spot of kill mud was set, before pulling out of the hole to 1994 m DRF began, ending at 1145 h. The BHA reached the surface at 2300 h, after which the tools were laid down and preparation for coring began. Transponders were released from the seafloor and all were recovered by the Hakuryu-Maru by 1500 h on 7 December. The vessel then began to move to Site C0002 for coring operations.

Hole C0018B

The Chikyu moved to Site C0018 (water depth = 3084.5 mbsl) after completing all operations at Site C0002. Making up and running the LWD BHA started at 1400 h on 25 December. The 12¼ inch LWD BHA components were similar to Hole C0012H operations with the exception of the sonicVISON tool. Function tests on the LWD were conducted at 76, 1060, 2006, and 3100 m DRF, respectively, while lowering the LWD BHA. Spud-in of Hole C0018B was confirmed by real-time resistivity data at 0430 h on 26 December. Washing down with a controlled ROP at an average of 40 m/h (maximum = 50 m/h) changed to drilling ahead from 40 mbsf at 0600 h. Rotation speed was gradually increased, step-wise, to TD (350 mbsf): 15 rotations per minute (rpm) × 1.8–2.5 kNm (40–45 mbsf), 30 rpm × 0–5.0 kNm (45–50 mbsf), 45 rpm × 0–5.0 kNm (50–67 mbsf), 60 rpm × 0–5.0 kNm (67–242 mbsf), and 80 rpm × 1.6–8.3 kNm (242–350 mbsf). The target depth was reached by 1130 h on 26 December. After spotting kill mud, the LWD BHA was pulled out of the hole to 2813 m DRF (300 m above the seafloor) by 2000 h on 26 December in preparation for operations at Site C0021.

Hole C0021A

The Chikyu shifted from Hole C0018B, where LWD/MWD operations had just been completed, and prepared to begin running the LWD drilling BHA into the hole from 2200 h on 26 December. The seabed was tagged (water depth = 2940.5 mbsl) at 2330 h, and washing down to 40 mbsf in Hole C0021A commenced. From 40 mbsf, drilling ahead began at 0100 h on 27 December, again increasing rotation speed in gradual stepped increases to TD (294.0 mbsf): 15 rpm × 1.8–2.5 kNm (40–45 mbsf), 30 rpm × 0–5.0 kNm (45–50 mbsf), 45 rpm × 0–5.0 kNm (50–60 mbsf), 60 rpm × 0–5.0 kNm (67–103 mbsf), and 80 rpm × 1.6–8.3 kNm (103–294 mbsf). TD was reached at 1100 h, and the hole was spotted with kill mud. While the LWD BHA was pulled out of the hole to the surface, overpull (80 kN) was observed from 88 to 101 mbsf; no other overpull was observed. The tools were broken and racked back to the derrick while the vessel moved to the next site by 2030 h in preparation for LWD at Site C0022.

Site C0022

Hole C0022A

Operations began with dropping transponders at Site C0022 from 2030 h on 27 December. The transponders were all confirmed on the seabed by 0143 h on 28 December, whereupon DP calibration began, finishing at 0530 h. The vessel moved 1.5 nmi upcurrent from the well center and began making up and running the LWD BHA into the hole from 0730 h. The LWD BHA was run to 260 m DRF by 1445 h, when time was taken to test the active heave compensator (AHC). After 30 min in standby mode, the AHC test was suspended because the cause of the malfunction could not be found. Running into the hole continued to the seabed, when Hole C0022A was spudded at 1600 h (water depth = 2675.5 mbsl) and the next 40.5 mbsf was washed down. From 1730 h, LWD drilling commenced following the same rotation speed increase with depth plan to TD as for Holes C0018B and C0021A (15 rpm × 0.6–2.5 kNm [40.5–45.5 mbsf], 30 rpm × 0.9–3.4 kNm [45.5–50.5 mbsf], 45 rpm × 0.5–4.7 kNm [50.5–60.5 mbsf], 60 rpm × 0.8–4.4 kNm [60.5–82.5 mbsf], 70 rpm × 1.1–7.9 kNm [82.5–111.5 mbsf], and 80 rpm × 0.9–14.1 kNm [111.5–136 mbsf]). Kill mud was spotted from 2200 h, and the LWD BHA was pulled out of the hole to 36 mbsf for WOW. At 0300 h on 29 December, the LWD BHA was run down to 96 mbsf and then reamed down from 0330 h to 136.0 mbsf. Once back at 136.0 mbsf, drilling ahead with the LWD BHA resumed to TD (420.5 mbsf), finishing at 1645 h. From 1715 h the LWD BHA was pulled out of the hole to 2692.6 m DRF (11 m above seabed), while a damaged drill pipe joint was laid down. Pulling out of the hole continued from 1930 h, reaching the surface by 0315 h on 30 December.

Hole C0022B

At 1800 h on 30 December, the vessel moved 10 nmi north of Site C0022, into the LCA, to begin loading bentonite and drilling equipment. From 2130 h, the HPCS/ESCS BHA was made up and tested. From 0115 h the BHA was run into the hole and the vessel began drifting toward the well center by 0300 h on 31 December. The HPCS BHA reached 2960 m DRF by 0745 h and for the next 2 h stood by while the core BOP ball valve was worked on. The bit was lowered to 2700 m DRF (2 m above seabed), the sinker bar was run down at 1130 h in preparation for cutting the first core, and spud-in and shooting took place at 1245 h (seafloor = 2674 mbsl). Washing down to 19.5 mbsf ended at 1500 h, and piston coring resumed at 1600 h. By 0515 h on 1 January, HPCS Core 338-C0022B-8H had been cut at 84.5 mbsf. At that point, coring operations switched from HPCS to EPCS, cutting the first EPCS core (9T) by 0945 h to 89.5 mbsf. At 1730 h, the coring system was changed from EPCS to ESCS, cutting from 99.5 mbsf. ESCS coring continued, pausing at 1315 h on 2 January to drill ahead from 171.5 to 190.5 mbsf. ESCS coring from 190.5 mbsf resumed at 1715 h. There was another pause in ESCS coring on 0315 h on 3 January, while drilling ahead from 228.5 to 266.5 mbsf (38 m advance) was completed. ESCS Core 338-C0022B-24X was cut from 0900 h, after which coring resumed. Coring in Hole C0022B finished at 0745 h on 5 January with the recovery of Core 41X. Kill mud was spotted, and the ESCS BHA was pulled out of the hole to the surface by 2330 h. As drill pipe was laid down, the transponders were released and recovered on the surface by the watch boat. The vessel began moving to Site C0021 in preparation for coring operations.

Hole C0021B

After coring operations were completed in Hole C0022B, preparations began for testing the small-diameter rotary core barrel (SD-RCB) system at 1515 h on 6 January; however, during make-up of the SD-RCB BHA, it was found that the head sub would not fit the other SD-RCB parts, so the test was cancelled at 1900 h. From that time, the HPCS/ESCS coring BHA making up and running into the hole began for coring at Site C0021. By 0300 h on 7 January, the BHA had reached 2930.7 m DRF while drifting in to Site C0021 (the well center was reached by 0145 h). From 0330 h, troubleshooting the crown-mounted heave compensator (CMC) began, during which operations continued in stand-by while the main HPS line was also cut and slip. Upon completion at 1030 h, the coring BHA continued to be run into the hole to 2946 m DRF. The coring BHA was set to capture the first core while still above the seafloor, so at 1145 h, while the BHA was at 2969 m DRF, the HPCS inner barrel was shot, penetrating 5.9 m. The seafloor was then determined as being 2972.5 m DRF (2944.0 mbsl). After the first core was taken, the BHA was washed down to 80 mbsf before coring resumed at 1515 h. There was a short pause in coring from 1845 to 2045 h while the wireline BOP hydraulic hose leak was fixed. Coring resumed from Core 338-C0021B-5H at 108.5 mbsf from 2100 h, and ended at Core 12H; Cores 10H, 11H, and 12H ended up becoming stuck in the inner core barrel to greater and lesser degrees; therefore, coring continued with the EPCS for the last two cores. Efforts to recover the stuck core liners continued on the rig floor until 9 January as coring continued. EPCS Core 338-C0021B-13T was cut at 1730 h from 175.5 mbsf, and coring finished with EPCS Core 14T to a final TD of 194.5 mbsf at 2100 h on 8 December. Kill mud was spotted at 2230 h and the EPCS BHA was pulled out of the hole to the surface, finishing at 0130 h on 10 January, ending operations related to Site C0021. The ship began transit to Shimizu port, Japan, reaching port on the morning of 11 January 2013. The scientists disembarked on 12 January 2013; the official end of the expedition was 13 January.